### v2

# Large recurrence of the polynomial Van der Waerden’s theorem and its application in Ramsey theory

Abstract
Recently M. Di Nasso and M. Ragosta proved an exponential version of Hindman’s finite sum theorem: for every finite partition of $\mathbb{N}$, there exists a monochromatic infinite exponential pattern of a single sequence. They used the stronger version of Furstenberg’s Central Sets Theorem and the theory of ultrafilters to prove their results. In this article, we prove these patterns are abundant by considering two-fold extensions of the Di Nasso-Ragosta theorem. We prove that each term of this pattern can be chosen from a finite set of any given length, consisting of a multiplicative $IP_{r}$ (that we will define later) set and a set consisting of geometric progression (sets containing both arithmetic and geometric progression) of finite length. To prove our result, first, we improve the size of the recurrence set of V. Bergelson and A. Leibmann’s polynomial Van der Waerden’s Theorem, and then we study its two-fold applications. First, we study monochromatic patterns involving polynomials over sum sets in additive piecewise syndetic sets. Then in the second, applying methods of induction, we construct infinitary patterns involving a two-fold improvement of the Di Nasso-Ragosta theorem. Then we study the properties of some special classes of ultrafilters witnessing variants of these patterns. Throughout our article, we use the methods from the Algebra of the Stone-Čech-Compactification of discrete semigroups (for details see [31]).
Keywords: Polynomial van der Waerden theorem, Exponential Hindman theorem, Piecewise syndetic set, Central set, ultrafilters, Algebra of the Stone-Čech compactification of discrete semigroup
Mathematics subject classification 2020: 05D10, 05C55, 22A15, 54D35
1   Introduction
Arithmetic Ramsey theory deals with the monochromatic patterns found in any given finite coloring of the integers or of the natural numbers $\mathbb{N}$. Here, “coloring” means disjoint partition, and a set is called “monochromatic” if it is included in one piece of the partition. Let $\mathcal{F}$ be a family of finite subsets of $\mathbb{N}.$ If for every finite coloring of $\mathbb{N}$, there exists a monochromatic member of $\mathcal{F}$, then such a family $\mathcal{F}$ is called a partition regular family. So basically, Ramsey theory is the study of the classifications of partitioned regular families. Arguably the first substantial development in this area of research was due to I. Schur [37] in $1916,$ when he proved that the family $\{\{x,y,x+y\}:x\neq y\}$ is a partitioned regular family over $\mathbb{N}.$ The second cornerstone development was due to Van der Waerden in $1927$ when he proved the following theorem.
Theorem 1.1 (Van der Waerden theorem, [40]).
For any finite coloring of the natural numbers one always finds arbitrarily long monochromatic arithmetic progressions. In other words, the set of all arithmetic progressions of finite length is partitioned regular.
Nonlinear versions of classical Ramsey theoretic are difficult to prove. One of the first progress in this direction was due to V. Bergelson and A. Liebman in [12], where they proved the nonlinear version of van der Waerden’s theorem, known as Polynomial van der Waerden’s Theorem. The authors proved a generalized version of the much stronger Szémeredi’s Theorem, but we are not going to discuss it in this paper. They used the methods from the topological dynamics and the PET induction, introduced in [3]. Throughout this article, we assume $\mathbb{P}$ to be the set of all polynomials from $\mathbb{Z}$ to $\mathbb{Z}$ with no constant term.
Theorem 1.2 (Polynomial van der Waerden Theorem, [12]).
Let $r\in\mathbb{N},$ and $\mathbb{N}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring of $\mathbb{N}$. Then for any finite collection of polynomials $F$ in $\mathbb{P}$, there exist $a,d\in\mathbb{N}$ and $1\leq j\leq r$ such that $\{a+p(d):p\in F\}\subset C_{j}.$
We will come back to our discussion after a short while.
Hindman Theorem and its variant
After Schur proved his result, an immediate question appeared: if there exists any infinitary extension of Schur’s theorem. This was a longstanding conjecture until $1974,$ when N. Hindman [24] solved this conjecture. Before we state his theorem, we need some technical definitions.
Definition 1.3 ($IP$-set).
Let $(S,+)$ be any commutative semigroup. Define
  1. 1.
    for any nonempty set $X,$ let $\mathcal{P}_{f}(X)=\{A\subseteq X:|A|<\infty\},$ and
  2. 2.
    for any injective sequence $\langle x_{n}\rangle_{n},$ define $\operatorname{FS}(\langle x_{n}\rangle_{n})=\left\{\sum_{t\in H}x_{t}:H\in\mathcal{P}_{f}(\mathbb{N})\right\},$
  3. 3.
    for any injective sequence $\langle x_{n}\rangle_{n},$ and for $\alpha\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ we write $x_{\alpha}=\sum_{n\in\alpha}x_{n}$,
  4. 4.
    a set $A\subseteq S$ is said to be an $IP$ set if there exists an injective sequence $\langle x_{n}\rangle_{n},$ such that $A=\operatorname{FS}(\langle x_{n}\rangle_{n}),$
  5. 5.
    for any $r\in\mathbb{N}$, a set $A\subseteq S$ is said to be an $IP_{r}$ set if there exists a sequence $\langle x_{n}\rangle_{n=1}^{r}$ such that $A=\operatorname{FS}(\langle x_{n}\rangle_{n=1}^{r})=\{\sum_{t\in H}x_{t}:(\neq\emptyset)H\subseteq\{1,2,\ldots,r\}\}.$
In our work, we use the terminology $IP/IP_{r}$ set to denote the corresponding set in $(\mathbb{N},+)$. Similarly, the terminology $MIP/MIP_{r}$ is set to denote the corresponding set in $(\mathbb{N},\cdot)$. The following theorem is known as Hindman theorem/ Finite sum theorem.
Theorem 1.4 (Hindman Theorem, [24]).
For every finite coloring of $\mathbb{N}$, there exists a monochromatic $IP$ set.
Recent research has focused on finding exponential patterns in the Ramsey theory. In [39], A. Sisto first initiated such kinds of works. Using the ultrafilter approach, he first proved the existence of an exponential Schur triple (i.e., patterns of the form $\{x,y,x^{y}:x\neq y\}$) for two colorings of the set of naturals. Later, using color-focusing arguments, J. Sahasrabudhe [35] proved this result for finite colorings. Later in [31], using tensor products of ultrafilters, M. Di Nasso and M. Ragosta proved a relatively strong version of this result. Then, in [18], they proved an exponential version of the Hindman theorem. Basically, they proved a key theorem [18, Theorem 3.1], and as a consequence, we have the following exponential version of Hindman’s theorem.
Theorem 1.5 (Exponential Hindman theorem, [18]).
For every finite partition of $\mathbb{N}$, there exists an injective sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ such that
$FEP(\langle x_{n}\rangle_{n})=\left\{x_{i_{n}}^{x_{i_{n-1}}^{\cdot^{\cdot^{\cdot^{x_{i_{1}}}}}}}:1\leq i_{1}<\cdots<i_{n-1}<i_{n},n\in\mathbb{N}\right\}$
is monochromatic.
We refer to such a pattern as Hindman Tower. Before we proceed further, we pause our attention to recall basic facts from the algebra of the Stone-Čech compactification of discrete semigroups.
Ultrafilters and its connections with Ramsey theory
Ultrafilters are set-theoretic objects that are intimately related to Ramsey theory. In this section, we give a brief introduction to this theory. For details, we refer to the book [31] to the readers. For a quick review authors can see the article [5]. A filter $\mathcal{F}$ over any nonempty set $X$ is a collection of subsets of $X$ such that
  1. 1.
    $\emptyset\notin\mathcal{F}$, and $X\in\mathcal{F}$,
  2. 2.
    $A\in\mathcal{F}$, and $A\subseteq B$ implies $B\in\mathcal{F},$ and
  3. 3.
    $A,B\in\mathcal{F}$ implies $A\cap B\in\mathcal{F}.$
Using Zorn’s lemma we can guarantee the existence of maximal filters which are called ultrafilters. Any ultrafilter $p$ has the following property:
  • if $X=\bigcup_{i=1}^{r}A_{i}$ is any finite partition of $X$, then there exists $i\in\{1,2,\ldots,r\}$ such that $A_{i}\in p.$
Therefore to prove a certain structure is monochromatic, it is sufficient to study ultrafilters each of whose members contain that structure.
Space of ultrafilters
Let $(S,\cdot)$ be a discrete semigroup. Denote by $\beta S$ the collection of all ultrafilters over $S.$ For any $(\emptyset\neq)A\subseteq S,$ let $\overline{A}=\{p\in\beta S:A\in p\}.$ The collection of sets $\{\overline{A}:(\emptyset\neq)A\subseteq S\}$ forms a basis for a topology. One can show with this topology $\beta S$ is the Stone-Čech compactification of $S.$ For any $A\subseteq S,$ and $x\in S$ denote $x^{-1}A=\{y:x\cdot y\in A\}.$ Now one can naturally extend the operation $\cdot$’ over the set $\beta S:$ for any $p,q\in\beta S$, let us define $p\cdot q$ by
$A\in p\cdot q\iff\{x:x^{-1}A\in q\}\in p.$
One can show with this definition $(\beta S,\cdot)$ is a semigroup. In fact one can prove that $(\beta S,\cdot)$ is a compact right topological semigroup. Here right topological means right action is continuous: for each $q\in\beta S,$ the map $\rho_{q}:\beta S\rightarrow\beta S$ defined by $\rho_{q}(p)=p\cdot q$ is continuous. However, the left action is not always continuous but: for $x\in S,$ the map $\lambda_{x}:\beta S\rightarrow\beta S$ defined by $\lambda_{x}(p)=x\cdot p$ is continuous. Now from Ellis [19] Theorem, we can say that any closed subsemigroup of $(\beta S,\cdot)$ contains idempotents. However, with some little modification, we can define the notions of $IP$ sets over any arbitrary semigroup, but we are not going for that. One can see that
  • If $(S,\cdot)$ is a semigroup, then a set $A\subseteq S$ contains an $IP$ set if and only if there exists an element $p\in E(\beta S,\cdot)$ such that $A\in p.$
Note that if $f:(S,\cdot)\rightarrow(T,\cdot)$ be any map between discrete semigroups. Then $\tilde{f}:\beta S\rightarrow\beta T$ be the continuous extension of $f$ defined by $\tilde{f}(p)=\tilde{f}(\lim_{x\rightarrow p}x).$
Notions of large sets
In the study of ultrafilter theory, one can show some members of $\beta S$ are combinatorial rich. In fact one can characterize combinatorially the members of such ultrafilters. During our study we will apply both combinatorial and algebraic characterizations of those ultrafilters. Here we summarize some of the notions of large sets intimately related to our study. For detailed study on large sets, we refer to the articles [6, 27, 28].
Definition 1.6.
Let $(S,\cdot)$ be a semigroup, let $n\in\mathbb{N}$ and let $A\subseteq S$. We say that
  1. 1.
    for any subset $T\subseteq(\beta S,\cdot)$, we let $E(T)$ be the collection of all idempotents of $T;$
  2. 2.
    $A$ is a thick set if for any finite subset $F\subset S$, there exists an element $x\in S$ such that $Fx=\{fx:f\in F\}\subset A$;
  3. 3.
    $A$ is a syndetic set if there exists a finite set $F\subset S$ such that $S=\bigcup_{x\in F}x^{-1}A$, where $x^{-1}A=\{y:xy\in A\}$;
  4. 4.
    $A$ is piecewise syndetic set if there exists a finite set $F\subset S$ such that $\bigcup_{x\in F}x^{-1}A$ is a thick set. It is well known that $A$ is piecewise syndetic if and only if there exists $p\in K(\beta S,\cdot)$ such that $A\in p.$
A set $L\subseteq\beta S$ is said to be a left ideal if $\beta S\cdot L=\{q\cdot p:p\in L,q\in\beta S\}\subseteq L.$ Similarly $R\subseteq\beta S$ is said to a left ideal if $R\cdot\beta S=\{p\cdot q:p\in R,q\in\beta S\}\subseteq R.$ Again using Zorn’s lemma one can show that minimal left/ right ideals (w.r.t. inclusion) exist. It can be proved that minimal left ideals are closed. One can show that if $A\subseteq S$ is a piecewise syndetic set, then there exists a minimal left ideal $L$ of $(\beta S,\cdot)$, and $p\in L$ such that $A\in p.$ Denote by $K(\beta S,\cdot)$, the union of all minimal left ideals of $(\beta S,\cdot)$. It can be proved that $K(\beta S,\cdot)$ is also the union of all minimal right ideals of $(\beta S,\cdot)$. That is
$K(\beta S)$ =$=\bigcup\{L:L\text{ is a minimal left ideal of }\beta S\}$
=$=\bigcup\{R:R\text{ is a minimal right ideal of }\beta S\}.$
As each minimal left ideal is the image of $\beta S$ right action map, it is closed, and hence it contains idempotents. Hence, $E\left(K(\beta S,\cdot)\right)\neq\emptyset$.
Central sets play a vital role in Ramsey theory. The notion of central sets was first introduced by H. Furstenberg [20], using the methods of topological dynamics. Later in [7], V. Bergelson and N. Hindman first proved an equivalent version of central sets in $\mathbb{N}$, using ultrafilter theory, and then in [38], H. Shi and H. Yang extended this result for arbitrary semigroups. For details, on the history of Central sets, and its combinatorial applications, we refer to the articles [4, 26]. Here we recall the algebraic definition.
Definition 1.7 (Central set).
A set $A\subseteq S$ is said to be central if there exists $p\in E\left(K(\beta S,\cdot)\right)$ such that $A\in p.$
Thick sets and syndetic sets can also be characterized in terms of ultrafilters. It can be proved that a set $A\subseteq S$ is thick if and only if there exists a left ideal $L$ such that $L\subseteq\overline{A}$. And a set $A\subseteq S$ is syndetic if and only if for every left ideal $L$, $L\cap\overline{A}\neq\emptyset.$ Throughout our article, we follow the following convention:
  • A set $A\subseteq\mathbb{N}$ is additively (resp. multiplicatively) “large” if this set is large in $(\mathbb{N},+)$ (resp. in $(\mathbb{N},\cdot)$).
In our work, we consider a special class of family: “Linear Rado family”.
Definition 1.8 (Linear Rado family).
Let $\mathcal{F}$ be a family of finite subsets of $\mathbb{N}$. We call such family a linear Rado family if
  • (1)
    for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$, $n\cdot F\in\mathcal{F}$, and
  • (2)
    for any finite coloring of $\mathbb{N}$ there exists a monochromatic element of the form $F\in\mathcal{F}$,
Using the color compactness argument one can show that if $\mathcal{F}$ is a linear Rado family, then every multiplicatively piecewise syndetic set contains an element of the form $F\in\mathcal{F}$.
  • Hence for every $r\in\mathbb{N},$ every multiplicatively piecewise syndetic set contains a set of the form $\{d,a+id:i\in\{0,1,\ldots,r\}\}$, and an $IP_{r}$ set.
If $\mathcal{F}$ is a family of sets, then a set $A$ is dual to this family if $A\cap F\neq\emptyset$ for every $F\in\mathcal{F}$. After we prove a certain structure is preserved in a large set, an immediate question appears: how much structure is preserved in a single cell? In this type of study, we need these notions of dual sets. Some well-known dual sets that we use are $IP^{\star}$ sets, $IP_{r}^{\star}$ sets, etc. Note that if $\mathcal{F}$ and $\mathcal{G}$ are two families of subsets of $S$ satisfying $\mathcal{F}\subseteq\mathcal{G}$, then each $\mathcal{G}^{\star}$ set is also a $\mathcal{F}^{\star}$. As a consequence, every $IP_{r}^{\star}$ set is an $IP^{\star}$ set.
In [12], V. Bergelson and A. Leibman proved a relatively stronger version of the polynomial van der Waerden theorem. Later, M. Walter [41] found a simple combinatorial proof that uses color-focusing arguments. Then, in [25], N. Hindman found a relatively easy algebraic proof. For some recent developments on polynomial progressions we refer to the articles [9, 10, 14]. Here we address the following stronger version of the Polynomial van der Waerden theorem.
Theorem 1.9 (Stronger polynomial van der Waerden theorem, [12, 25]).
Let $A\subseteq\mathbb{Z}$ be a piecewise syndetic set, and $F\in\mathcal{P}_{f}(\mathbb{P})$. Then
$R=\left\{n:\left\{m:\left\{m,m+p\left(n\right):p\in F\right\}\subset A\right\}\text{ is piecewise syndetic }\right\}\text{}$
is an $IP^{\star}$ set.
Structure of the paper
The primary goal of this paper is to study how abundant the Hindman towers are. To achieve our goal, first, in Section 2.2, we improve Theorem 1.9. Under the same hypothesis, we prove $R$ is an $IP_{r}^{\star}$ set for some $r\in\mathbb{N}$ (see Theorem 2.3). The key tool to prove this result is the polynomial Hales-Jewett Theorem, originally proved by V. Bergelson and N. Hindman [41]. As a corollary, we see that exponential Schur triples are abundant. In Section 2.2.1, we prove additive piecewise syndetic sets contain more rich patterns involving polynomials over “sum subsystems”. Introducing a new induction technique in [22, Theorem 4.1.], the author of this paper, in collaboration with L. L. Bagilini and S. K. Patra, found the polynomial extension of the key theorem of Di Nasso and Ragosta (Theorem 3.1).  In Section 3, we improve the Di Nasso-Ragosta Theorem [18, Theorem 3.1] in two different directions. As a consequence, we show the existence of monochromatic Hindman’s towers, where each element of these towers comes from a set containing an $MIP_{r}$ set (for any given $r\in\mathbb{N}$), and a set of geometric progressions of length $r.$ We also study the properties of these ultrafilters, witnessing the aforementioned patterns.
2   Large recurrence of the polynomial Van der Waerden’s theorem
Hales-Jewett theorem [23] is one of the fundamental pillars of Ramsey theory. Many classical results in Ramsey theory directly followed from this theorem. The strength of this theorem is that the statement of this theorem is set-theoretical, and different algebraic operations induce different Ramsey theoretic statements. For example, van der Waerden’s theorem is an immediate consequence of this theorem. The original proof of the Hales-Jewett theorem uses simple combinatorial arguments. Later people found dynamical and algebraic proofs. One may see [4, 32] for reference. Soon after V. Bergelson and A. Leibman proved the polynomial van der Waerden theorem (Theorem 1.2), an immediate question appears if there exists any polynomial extension of the Hales-Jewett theorem. Later in [13], V. Bergelson and A. Leibman established the Polynomial extension of the Hales-Jewett theorem. The proof uses methods from Topological dynamics and PET induction. Then in [41], M. Walter found a combinatorial proof. In [29, 30], N. Hindman and R. McCutcheon extensively studied this theorem using the algebraic approach. Using the Polynomial Hales-Jewett theorem, in this section, first we refine Theorem 1.9. Then we use it to deduce some new results.
Revisiting Polynomial Hales-Jewett Theorem
Now we pause to recall the Hales-Jewett Theorem and it’s polynomial extension. Let $\omega=\mathbb{N}\cup\left\{0\right\}$, where $\mathbb{N}$ is the set of positive integers. Given a nonempty set $\mathbb{A}$ called alphabet, a finite word is an expression of the form $w=a_{1}a_{2}\ldots a_{n}$ with $n\geq 1$ and $a_{i}\in\mathbb{A}$. The quantity $n$ is called the length of $w$ and denoted $\left|w\right|$. Let $v$ (a variable) be a letter not belonging to $\mathbb{A}$. By a variable word over $\mathbb{A}$ we mean a word $w$ over $\mathbb{A}\cup\left\{v\right\}$ that has at least one occurrence of $v$. For any variable word $w$, $w\left(a\right)$ is the result of replacing each occurrence of $v$ by $a$.
The following theorem is known as Hales-Jewett theorem, is due to A. W. Hales and R. I. Jewett.
Theorem 2.1.
[23, Hales-Jewett Theorem (1963)] For all values $t,r\in\mathbb{N}$, there exists a number $\text{HJ}\left(r,t\right)$ such that, if $N\geq\text{HJ}\left(r,t\right)$ and $\left[t\right]^{N}$ is $r$ colored then there exists a variable word $w$ such that $\left\{w\left(a\right):a\in\left[t\right]\right\}$ is monochromatic.
The word space $\left[t\right]^{N}$ is called Hales-Jewett space or H-J space. The number $\text{HJ}\left(r,t\right)$ is called Hales-Jewett number.
Before we recall the polynomial Hales-Jewett theorem we need some new notions. For $q,N\in\mathbb{N}$, let $Q=[q]^{N}$. For $a\in Q$, $\emptyset\neq\gamma\subseteq[N]$ and $1\leq x\leq q$, $a\oplus x\gamma$ is defined to be the vector $b$ in $Q$ obtained by setting $b_{i}=x$ if $i\in\gamma$ and $b_{i}=a_{i}$ otherwise.
In the statement of Theorem [41, Polynomial Hales-Jewett Theorem], we have $a\in Q$ so that $a=\langle\vec{a}_{1},\vec{a}_{2},\ldots,\vec{a}_{d}\rangle$ where for $j\in\{1,2,\ldots d\}$, $\vec{a}_{j}\in[q]^{N^{j}}$ and we have $\gamma\subseteq[N]=\{1,2,\ldots,N\}$. Given $j\in\{1,2,\ldots,d\}$, let $\vec{a}_{j}=\langle a_{j,\vec{i}}\rangle_{\vec{i}\in N^{j}}$. Then $a\oplus x_{1}\gamma\oplus x_{2}(\gamma\times\gamma)\oplus\ldots\oplus x_{d}\gamma^{d}=b$ where $b=\langle\vec{b}_{1},\vec{b}_{2},\ldots,\vec{b}_{d}\rangle$ and for $j\in\{1,2,\ldots,d\}$, $\vec{b}_{j}=\langle b_{j,\vec{i}}\rangle_{\vec{i}\in N^{i}}$ where
$b_{j,\vec{i}}=\left\{\begin{array}[]{rl}x_{j}&\hbox{if}\,\vec{i}\in\gamma^{i}\\ a_{j,\vec{i}}&\hbox{otherwise.}\end{array}\right.$
The following theorem is the Polynomial Hales-Jewett Theorem.
Theorem 2.2.
[41, Polynomial Hales-Jewett Theorem] For any $q,k,d$ there exists $N\left(q,k,d\right)\in\mathbb{N}$ such that whenever $Q=Q\left(N\right)=\left[q\right]^{N}\times\left[q\right]^{N\times N}\times\cdots\times\left[q\right]^{N^{d}}$ is $k$-colored there exist $a\in Q$ and $\gamma\subseteq\left[N\right]$ such that the set of points
$\left\{a\oplus x_{1}\gamma\oplus x_{2}\left(\gamma\times\gamma\right)\oplus\cdots\oplus x_{d}\gamma^{d}:1\leq x_{i}\leq q\right\}$
is monochromatic.
$IP_{r}^{\star}$ Polynomial van der Waerden’s theorem
As promised above, in this section we strengthen Theorem 1.9. We provide both algebraic and combinatorial proofs. Both of these proofs have their own pros and cons. The Algebraic proof is short but uses Zorn’s lemma, whereas the combinatorial proof is lengthy, a little complicated, but does not use Zorn’s lemma.
Theorem 2.3.
Let $A\subseteq\mathbb{N}$ be an additive piecewise syndetic set and $F\in\mathcal{P}_{f}(\mathbb{P})$. Then
$\left\{n:\left\{m:\left\{m,m+p\left(n\right):p\in F\right\}\subset A\right\}\text{ is piecewise syndetic }\right\}$
is an $IP_{r}^{\star}$ set for some $r\in\mathbb{N}.$
Proof.
(Combinatorial proof:) As $A$ is a piecewise syndetic set, there exists a finite set $F_{1}\subset\mathbb{N}$ such that $B=\cup_{t\in F_{1}}\left(-t+A\right)$ is a thick set. Assume that $|F_{1}|=k.$ Let $\mathbb{A}$ be the set of all coefficients of the polynomials of $F,$ and $d$ be the maximum degree of the polynomials of $F.$
To prove the given set is an $IP_{r}^{\star}$ set for some $r\in\mathbb{N}$, we need to show that the given set intersects with every set of the form $\operatorname{FS}(\langle x_{n}\rangle_{n=1}^{r})$ for some $r\in\mathbb{N}$. Our claim is that $r=N.$ Let $N=N\left(|\mathbb{A}|,k,d\right)\in\mathbb{N}$ be the polynomial Hales-Jewett number guaranteed by Theorem 2.2. Let $\langle x_{n}\rangle_{n=1}^{N}$ be any sequence in $\mathbb{N}.$ Let $Q=Q\left(N\right)$ be the set as in Theorem 2.2. Define the map $\gamma:Q\rightarrow\mathbb{Z}$ by
$\gamma\left(\left(a_{1},\ldots,a_{N},a_{11},\ldots,a_{NN},\ldots,a_{11\cdots 1(d\text{ times})},\ldots,a_{NN\cdots N(d\text{ times})}\right)\right)=$
$\sum_{i=1}^{n}a_{i}x_{i}+\sum_{i,j=1,1}^{N,N}a_{ij}x_{i}x_{j}+\cdots+\sum_{i_{1},i_{2},\ldots,i_{d}=1,1,\cdots,1}^{N,N,\cdots,N}a_{i_{1}i_{2}\cdots i_{d}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{d}}.$
If necessary, translate the set $G=\left\{\gamma\left(\overrightarrow{a}\right):\overrightarrow{a}\in Q\right\}$ by a sufficiently large number $M$ to assume that $G\subset\mathbb{N}$. Then for any finite coloring of $G,$ there exists $a\in\mathbb{N},$ and $d\in FS\left(\langle x_{n}\rangle_{n=1}^{N}\right)$ such that $\left\{a+p\left(d\right):p\in F\right\}\cup\left\{a\right\}$ is monochromatic.
As $G$ is finite, the number of such patterns in $G$ is finite. Let $\mathscr{H}$ be the collection of such patterns in $G$. Let $|\mathscr{H}|=l$, =$\mathscr{H}=\left\{H_{i}:1\leq i\leq l\right\}$ and $F_{1}=\{f_{i}:1\leq i\leq k\}.$ As $G$ is finite, the collection $C=\left\{x:G+x\subset B\right\}$ is a thick set.
Now construct a $[l]\times[k]$ coloring of $C$ by $x\in C$ has color $\left(i,j\right)\in[l]\times[k]$ if and only if there exists $f_{j}\in F_{1}$ and $H_{i}\in\mathscr{H}$ such that $\left\{x+f_{j}+m:m\in H_{i}\right\}\subset A$ (one can choose the least $\left(i,j\right)$ in dictionary ordering). Now $x\in C_{\left(i,j\right)}$ if and only if $x\in C$ and $x$ has color $\left(i,j\right).$ So, $C=\cup_{\left(i,j\right)\in[l]\times[k]}C_{\left(i,j\right)}.$ Hence at least one of $C_{(i,j)}^{\prime}s$, say $C_{\left(s,t\right)}$ is a piecewise syndetic set. Hence there exists $H=\left\{q+p\left(r\right):p\in F\right\}\cup\left\{q\right\}\in\mathcal{H}$ such that for all $x\in C_{\left(s,t\right)}$
$\left\{x+\left(t+q\right)+p\left(r\right):p\in P\right\}\cup\left\{x+\left(t+q\right)\right\}\subset A.$
As $C_{\left(s,t\right)}$ is a piecewise syndetic set, and as piecewise syndetic sets are translation invariant, we have $C_{\left(s,t\right)}+\left(t+q\right)$ is a piecewise syndetic set. This implies for all $m\in C_{\left(s,t\right)}+\left(t+q\right),$ we have
$\left\{m\right\}\cup\left\{m+p\left(d\right):p\in F\right\}\subset A,$
for some $d\in FS\left(\langle x_{n}\rangle_{n=1}^{N}\right).$ Hence
$\left\{n:\left\{m:\left\{m,m+p\left(n\right):p\in F\right\}\subset A\right\}\text{ is piecewise syndetic }\right\}$
is an $IP_{N}^{\star}$. This completes the proof. ∎
The key idea of the following algebraic proof is to use [31, Theorem 4.39] to reduce the problem from piecewise syndetic sets to syndetic sets.
Proof.
(Algebraic proof:) Let $A$ be a piecewise syndetic set. Hence there exists $p\in K\left(\beta\mathbb{N},+\right)$ such that $A\in p.$ Hence by [31, Theorem 4.39], the set $S=\left\{x:-x+A\in p\right\}$ is a syndetic set. Therefore we have a finite set $F\subset\mathbb{N}$ such that $\mathbb{N}=\cup_{t\in F}-t+S.$ Assume that $|F|=r.$ Proceeding along the same line of the above proof, let us choose $N=N\left(\mathbb{A},r,d\right)\in\mathbb{N}$, a sequence $\langle x_{n}\rangle_{n=1}^{N}$ in $\mathbb{N}$, and the corresponding finite set $G\subset\mathbb{N}$. As $G\subset\cup_{t\in F}-t+S,$ there exists $(\emptyset\neq)H\subseteq\left\{1,2,\ldots,N\right\}$, and $b\in\mathbb{N}$ such that $\left\{b,b+f\left(x_{H}\right):f\in F\right\}\subset S.$ In other words
$D=\left\{d:\text{ there exists }m\text{ such that }\left\{m,m+f\left(d\right):f\in F\right\}\subset S\right\}$
is $IP_{N}^{\star}$ set. Now for each $n\in D,$ there exists $m$ such that $\left\{m,m+f\left(n\right):f\in F\right\}\subset S.$ Hence
$B=A\cap\left(-m+A\right)\cap\bigcap_{f\in F}-\left(m+f\left(n\right)\right)+A\in p.$
As $B$ is a piecewise syndetic set, $B+m$ is also a piecewise syndetic, and note that for each $a\in B+m,$ we have $\left\{a,a+f\left(n\right):f\in F\right\}\subset A.$ This completes the proof. ∎
It is easy to check that $\overline{E\left(K(\beta\mathbb{N},+)\right)}$ is a left ideal of $(\beta\mathbb{N},\cdot)$, and so it contains a minimal left ideal of $(\beta\mathbb{N},\cdot)$. Therefore we can choose $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{E\left(K(\beta\mathbb{N},+)\right)}.$ By choice of $p,$ each member of $p$ is additive and multiplicative central set. The following corollary immediately follows from Theorem 2.3.
Corollary 2.4.
Let $F\in\mathcal{P}_{f}(\mathbb{P})$, and $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{E\left(K(\beta\mathbb{N},+)\right)}.$ Then for every $A\in p,$
$\left\{d\in A:\left\{a:\left\{a,a+p\left(d\right):p\in F\right\}\subset A\right\}\text{ is additive piecewise syndetic }\right\}$
is both additive and multiplicative piecewise syndetic.
Proof.
Given $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{E\left(K(\beta\mathbb{N},+)\right)}.$ As $A\in p,$ we have $A$ is an additive and multiplicative piecewise syndetic set. Hence from Theorem 2.3 we have
$D=\left\{d:\left\{a:\left\{a,a+p\left(d\right):p\in F\right\}\subset A\right\}\text{ is additive piecewise syndetic }\right\}$
is an $IP_{r}^{\star}$ set for some $r\in\mathbb{N}.$ As for each $r\in\mathbb{N}$, every multiplicative piecewise syndetic set contains an $IP_{r}$ set, we have $\overline{K\left(\beta\mathbb{N},\cdot\right)}\subset\overline{D}.$ So, $D\in p$ and as $A\in p$, we have $A\cap D\in p.$ Hence we have our desired conclusion. ∎
Let $\mathbb{N}_{>1}=\mathbb{N}\setminus\{1\},$ and $n\in\mathbb{N}_{>1}.$ A set $A\subseteq\mathbb{N}$ will be called $n$-exponential piecewise syndetic if there exists a multiplicative piecewise syndetic set $B$ such that $n^{B}=\{n^{b}:b\in B\}\subseteq A.$ For any $p\in\beta\mathbb{N},$ denote by $E_{1}(n,p)\in\beta\mathbb{N}$ such that $A\in E_{1}(n,p)$ if and only if $B=\{m:n^{m}\in A\}\in p.$ Clearly $A$ is $n$-exponential piecewise syndetic if and only if there exists $p\in\overline{K(\beta\mathbb{N},\cdot)}$ such that $A\in E_{1}(n,p).$
Passing to the map $\overrightarrow{a}$ to $\gamma(\overrightarrow{a})$ in the proof of Theorem 2.3, one can prove the following theorem and so we leave it to the reader.
Theorem 2.5.
Let $n\in\mathbb{N}_{>1}$, $F\in\mathcal{P}_{f}(\mathbb{P})$ and $A\subseteq\mathbb{N}$ be an multiplicative piecewise syndetic set. Then there exists $r\in\mathbb{N}$ such that
$\left\{b:\left\{a:\left\{a,a\cdot n^{p\left(b\right)}:p\in F\right\}\subset A\right\}\text{ is multiplicative piecewise syndetic }\right\}$
is an $IP_{r}^{\star}$ set.
The following corollary of the above theorem shows that exponential Schur patterns are abundant.
Corollary 2.6.
Let $n\in\mathbb{N}_{>1}$, $p\in\overline{K(\beta\mathbb{N},\cdot)}$ and $A\in E_{1}(n,p)$. Then
$\{y\in A:\{x:\{x,x^{y}\}\subset A\}\text{ is a }n\text{-exponential piecewise syndetic}\}\in E_{1}(n,p).$
Proof.
Let $A\in E_{1}(n,p)$ and so $B=\{m:n^{m}\in A\}\in p.$ Let $P(x)=x$ be the identity polynomial on $\mathbb{N}$. Now from Theorem 2.5,
$C=\left\{b:\left\{a:\left\{a,a\cdot n^{b}\right\}\subset A\right\}\text{ is piecewise syndetic }\right\}$
is an $IP_{r}^{\star}$ set for some $r\in\mathbb{N}.$ As $C$ is an $IP_{r}^{\star}$ set, we have $D=A\cap C\in p.$ Hence
$D=\left\{b\in A:\left\{a:\left\{a,a\cdot n^{b}\right\}\subset A\right\}\text{ is piecewise syndetic }\right\}\in p.$
Now for every $m\in\mathbb{N}$, passing to the map $m\rightarrow n^{m}$ we have our desired result. ∎
Before we proceed to the next section, here we give a short proof of the existence of monochromatic $\{x,y,x^{y}\},$ which is not related to this work as much. Though we have a very short proof [22, Page 3: Direct proof of Theorem 1.5.], still we think that the following proof should deserve a place in the literature.
Let $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{E\left(K(\beta\mathbb{N},+)\right)},$ and $A\in p.$ Define a map $P:(\mathbb{N}\cup\{0\},+)\rightarrow(\mathbb{N},\cdot)$ by $P(n)=2^{n}.$ This map clearly satisfies the definition of the polynomial map between countable commutative semigroups [11, Definition 2.8.]. Now $p\in\overline{E\left(K(\beta\mathbb{N},+)\right)}$ implies that there exists a sequence $\langle x_{n}\rangle_{n}$ in $\mathbb{N}$ such that $\operatorname{FS}\left(\langle x_{n}\rangle_{n}\right)\subseteq A.$ Now from the [11, Theorem 4.10], there exists $a\in\mathbb{N}$, and $b\in\operatorname{FS}\left(\langle x_{n}\rangle_{n}\right)$ such that $\{a,aP(b)=a2^{b}\}\subset A.$ Again $b\in A$. Hence $2^{A}$ contains such patterns. This implies $E_{1}(2,p)$ witness this pattern. However we choose any $n\in\mathbb{N}_{>1}$ instead of $2$
In our proof we used a relatively weak version of the polynomial Central Sets Theorem, however, the stronger version can be found in [30, 21].
Applications to new polynomial patterns
In this section, we deduce two new applications of Theorem 2.3. The first one involves patterns in additively piecewise syndetic sets, where the domain of the polynomials can be chosen from any given multiplicative central set. And the second one involves the evaluation of a polynomial over the “sum subsystem” of any given $IP$ set.
Theorem 2.7.
Let $F\in\mathcal{P}_{f}(\mathbb{P})$, $A$ be an additively piecewise syndetic set, and $B$ be a multiplicatively central set. Then there exists a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ such that
  1. (1)
    $FP(\langle x_{n}\rangle_{n=1}^{\infty})\subset B$, and
  2. (2)
    for each $N\in\mathbb{N}$, there exists $a(N)\in\mathbb{N}$ such that
    $\left\{a(N),a(N)+P(y):P\in F,y\in FS(\langle x_{n}\rangle_{n=1}^{N})\cup FP(\langle x_{n}\rangle_{n=1}^{N})\right\}\subset A.$
In addition for each $N\in\mathbb{N}$, such collection of $a(N)$ is piecewise syndetic.
Proof.
Given $B$ is a multiplicatively central set. Hence there exists $q\in E(K(\beta\mathbb{N},\cdot))$ such that $B\in q$. By [31, Lemma 4.14], we have $B^{*}=\{n\in B:n^{-1}B\in q\}\in q$. As $A$ is additively piecewise syndetic set, from Theorem 2.3 we have
$D=\{d:\{a:\{a,a+P(d):P\in F\}\subset A\}\text{ is a piecewise syndetic}\}$
is an $IP_{r}^{*}$ for some $r\in\mathbb{N}$. Hence, $\overline{K(\beta\mathbb{N},\cdot)}\subset\overline{D}$, and this implies $B^{*}\cap D\in q$. Let $x_{1}\in B^{*}\cap D$. Then there exists $a(1)\in\mathbb{N}$ such that $\{a(1),a(1)+P(x_{1}):P\in F\}\subset A$ and such collection of $a(1)$ is piecewise syndetic. Inductively assume that for some $N\in\mathbb{N}$, there exists a finite sequence $\langle x_{n}\rangle_{n=1}^{N}$ such that
  1. 1.
    $FP(\langle x_{n}\rangle_{n=1}^{N})\subset B^{*}$, and
  2. 2.
    for $1\leq k\leq N$, $\exists$ $a(k)\in\mathbb{N}$ such that $B_{k}=\{a:\{a,a+P(y):P\in F,y\in FS(\langle x_{n}\rangle_{n=1}^{k})\cup FP(\langle x_{n}\rangle_{n=1}^{k})\}\subset A\}$ piecewise syndetic.
Now $B_{N}$ is a piecewise syndetic set. For each $y\in FP(\langle x_{n}\rangle_{n=1}^{N})$, $z\in FS(\langle x_{n}\rangle_{n=1}^{N})$, and $P\in F$, define new polynomials $P_{y},P^{z}\in\mathbb{P}$ by $P_{y}(n)=P(ny)$, and $P^{z}(n)=P(z+n)-P(z)$ $\forall n\in\mathbb{Z}$. Define
$G=F\cup\{P_{y}:P\in F,y\in FP(\langle x_{n}\rangle_{n=1}^{N})\}\cup\{P^{z}:P\in F,z\in FS(\langle x_{n}\rangle_{n=1}^{N})\},$
and so $G\in\mathcal{P}_{f}(\mathbb{P})$. Again from Theorem 2.3,
$D_{1}=\{d:\{b:\{b,b+f(d):f\in G\}\subset B_{N}\}\text{ is a piecewise syndeticset}\}$
is an $IP_{s}^{*}$ for some $s\in\mathbb{N}$. From [31, Lemma 4.14], we know that for each $y\in FP(\langle x_{n}\rangle_{n=1}^{N})$, $y^{-1}B^{*}\in q.$ Hence we have $E=D_{1}\cap\bigcap_{y\in FP(\langle x_{n}\rangle_{n=1}^{N})}y^{-1}B^{*}\cap B^{*}\in q$. Choose $x_{N+1}\in E$, and so $FP(\langle x_{n}\rangle_{n=1}^{N+1})\subset B^{*}$. Again $x_{N+1}\in D_{1}$ and so there exists $a(N+1)$ such that
$\left\{a(N+1),a(N+1)+f(x_{N+1}):f\in G\right\}\subset B_{N}.$
Again the collection of such $a(N+1)$ is piecewise syndetic. Now,
  1. 1.
    $a(N+1)\in B_{N}$ implies $a(N+1)\in A$, and for all $y\in FP(\langle x_{n}\rangle_{n=1}^{N})\cup FS(\langle x_{n}\rangle_{n=1}^{N})$, and $P\in F$ we have $a(N+1)+P(y)\in A$, (A)
  2. 2.
    for all $P\in F,y\in FP(\langle x_{n}\rangle_{n=1}^{N})$, $a(N+1)+P_{y}(x_{N+1})\in A$ implies $\forall y\in FP(\langle x_{n}\rangle_{n=1}^{N+1})\setminus FP(\langle x_{n}\rangle_{n=1}^{N})$, $a(N+1)+P(y)\in A$, and (B)
  3. 3.
    for all $P\in F,z\in FS(\langle x_{n}\rangle_{n=1}^{N})$, $a(N+1)+P^{z}(x_{N+1})\in A$ implies $\forall z\in FP(\langle x_{n}\rangle_{n=1}^{N+1})\setminus FP(\langle x_{n}\rangle_{n=1}^{N})$, $a(N+1)+P(z)\in A$. (C)
Combining (A), (B), and (C), for all $P\in F$, and $y\in FS(\langle x_{n}\rangle_{n=1}^{N+1})\cup FP(\langle x_{n}\rangle_{n=1}^{N+1})$, we have $a(N+1)+P(y)\in A$. As the collection of such $a(N+1)$ is piecewise syndetic, we proved the induction hypothesis, and this proves our result.
For any finite coloring of $\mathbb{N}$, the following special case of Theorem 2.7 says that we can choose the set $FP(\langle x_{n}\rangle_{n=1}^{\infty})$ from the same color as of $a(N)+P(y)$.
Corollary 2.8.
Let $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{\left(K(\beta\mathbb{N},+)\right)}$, and $A\in p.$ Then from Theorem 2.7, for any $F\in\mathcal{P}_{f}(\mathbb{P})$, there exists a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ such that $FP(\langle x_{n}\rangle_{n=1}^{\infty})\subset A$, and for each $N\in\mathbb{N}$, there exists $a(N)\in\mathbb{N}$ such that
$\{a(N),a(N)+P(y):P\in F,y\in FS(\langle x_{n}\rangle_{n=1}^{N})\cup FP(\langle x_{n}\rangle_{n=1}^{N})\}\subset A.$
In addition for each $N\in\mathbb{N}$, such collection of $a(N)$ is piecewise syndetic.
Sum subsystems of IP sets play important role in Ramsey theory. These systems appear in many Ramsey theoretic patterns to strengthen classical statements. Here we recall the definition of sum subsystems.
Definition 2.9.
(Sum subsystem) For any $IP$ set $FS\left(\langle x_{n}\rangle_{n}\right)$, a sum subsystem of $FS\left(\langle x_{n}\rangle_{n}\right)$ is a set of the form $FS\left(\langle y_{n}\rangle_{n}\right)$, where for each $n\in\mathbb{N},$ $y_{n}$ is defined as follows:
  1. 1.
    there exists a sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ such that for each $n\in\mathbb{N}$, $\max H_{n}<\min H_{n+1}$, and
  2. 2.
    $y_{n}=\sum_{t\in H_{n}}x_{t}$ for all $n\in\mathbb{N}$.
The following theorem is a variant of Theorem 2.7, where the polynomials are evaluated over sum-subsystems of a given $IP$ set. In the later section we will use this Theorem to prove the first improvement of [22, Theorem 4.1.].
Theorem 2.10.
Let $A$ be an additive piecewise syndetic set and $F\in\mathcal{P}_{f}(\mathbb{P})$. Then for every injective sequence $\langle x_{n}\rangle_{n}$ there exists a sum subsystem $FS\left(\langle y_{n}\rangle_{n}\right)$ of $FS\left(\langle x_{n}\rangle_{n}\right)$ such that for every $N\in\mathbb{N},$ there exists $a(N)\in\mathbb{N},$ such that
$\left\{a(N)+p\left(y\right):y\in FS\left(\langle y_{n}\rangle_{n=1}^{N}\right)\cup FP\left(\langle y_{n}\rangle_{n=1}^{N}\right)\text{and }p\in F\right\}\subset A.$
In addition for each $N\in\mathbb{N}$, such collection of $a(N)$ is piecewise syndetic.
Proof.
For any $v\in E\left(\beta\mathbb{N},+\right)$ and any $A\in v,$ let $A^{\star}=\{n\in A:-n+A\in v\}.$ As $v$ is an idempotent, we have $A^{\star}\in v$. Let $v$ be an idempotent such that $FS\left(\langle x_{n}\rangle_{n}\right)\in v,$ and so $FS\left(\langle x_{n}\rangle_{n}\right)^{\star}\in v.$ From Theorem 1.9, we have $y_{1}\in FS\left(\langle x_{n}\rangle_{n}\right)^{\star}$ such that the set $B_{1}=\left\{a:\left\{a,a+p(y_{1}):p\in F\right\}\subset A\right\}$ is piecewise syndetic. So there exists an element $a\in B_{1}$ such that $\left\{a,a+p(y_{1}):p\in F\right\}\subset A.$ Inductively assume that for some $M\in\mathbb{N}$,
  1. 1.
    $FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)\subset FS\left(\langle x_{n}\rangle_{n}\right)^{\star},$
  2. 2.
    For each $1\leq N\leq M,$ we have
    $\left\{a(N)+p\left(y\right):y\in FS\left(\langle y_{n}\rangle_{n=1}^{N}\right)\cup FP\left(\langle y_{n}\rangle_{n=1}^{N}\right)\text{and }p\in F\right\}\subset A,$
  3. 3.
    For each $1\leq N\leq M,$ the set
    $B_{N}=\left\{a:\left\{a,a+p(y):p\in F,y\in FS\left(\langle y_{n}\rangle_{n=1}^{N}\right)\cup FP\left(\langle y_{n}\rangle_{n=1}^{N}\right)\right\}\subset A\right\}\text{ is piecewise syndetic. }$.
For each $p\in F,$ $y\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right),$ and $z\in FP\left(\langle y_{n}\rangle_{n=1}^{M}\right),$ define new polynomials $p_{y},p^{\prime}_{z}:\mathbb{N}\rightarrow\mathbb{N},$ by $p_{y}\left(n\right)=p\left(n+y\right)-p\left(y\right)$ and $p^{\prime}_{z}\left(n\right)=p(zn).$ Choose a new collection of polynomials $G=F\cup\left\{p_{y}:p\in F,y\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)\right\}\cup\left\{p^{\prime}_{z}:p\in F,z\in FP\left(\langle y_{n}\rangle_{n=1}^{M}\right)\right\}.$ From Theorem 1.9,
$D=\left\{x:\left\{c:\left\{c,c+p(x):p\in G\right\}\subset B_{M}\right\}\text{ is piecewise syndetic }\right\}$
is an $IP^{\star}$ set. Note that if $A$ is an $IP^{*}$ set, and $n\in\mathbb{N},$ then $n^{-1}A$ is again an $IP^{*}$ set. Hence,
$D_{1}=D\cap\bigcap_{z\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)}z^{-1}D\cap FS\left(\langle x_{n}\rangle_{n}\right)^{\star}\cap\bigcap_{y\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)}\left(-y+FS\left(\langle x_{n}\rangle_{n}\right)^{\star}\right)\in v.$
Choose $y_{M+1}\in D_{1}$, and an element $c_{1}\in E=\left\{c:\left\{c,c+p(y_{M+1}):p\in G\right\}\subset B_{M}\right\}.$ Hence $FS\left(\langle y_{n}\rangle_{n=1}^{M+1}\right)\subset FS\left(\langle x_{n}\rangle_{n}\right)^{\star}.$ Now we complete the induction.
  • Case 1: if $z\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)\cup FP\left(\langle y_{n}\rangle_{n=1}^{M}\right)\cup\left\{y_{M+1}\right\}$, then for all $p\in F,$ $c_{1}+p\left(z\right)\in B_{M}\subset A,$ and
  • Case 2: if $z\in FS\left(\langle y_{n}\rangle_{n=1}^{M+1}\right)\setminus\left(FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)\cup\left\{y_{M+1}\right\}\right)$, then we have $z=s+y_{M+1}$, for some $s\in FS\left(\langle y_{n}\rangle_{n=1}^{M}\right)$.
    So we have
    $c_{1}+p_{s}\left(y_{M+1}\right)=c_{1}+p\left(s+y_{M+1}\right)-p\left(s\right)\in B_{M},$
    But then $c_{1}+p\left(s+y_{M+1}\right)-p\left(s\right)+p\left(s\right)=c_{1}+p\left(z\right)\in A.$
  • Case 3: if $z\in FP\left(\langle y_{n}\rangle_{n=1}^{M+1}\right)\setminus\left(FP\left(\langle y_{n}\rangle_{n=1}^{M}\right)\cup\left\{y_{M+1}\right\}\right)$, then we have and $z=ty_{M+1}$ for some $t\in FP\left(\langle y_{n}\rangle_{n=1}^{M}\right).$ Then
    $c_{1}+p^{\prime}_{t}\left(y_{M+1}\right)=c_{1}+p\left(ty_{M+1}\right)=c_{1}+p(z)\in B_{M}.$
Now rename $E=B_{M+1}$. Now for each $a\in B_{M+1}\subset A,$ we have
$\left\{a+p\left(z\right):p\in F\text{ and }z\in FS\left(\langle y_{n}\rangle_{n=1}^{M+1}\right)\cup FP\left(\langle y_{n}\rangle_{n=1}^{M+1}\right)\right\}\subset A,$
and the set $B_{M+1}$ is piecewise syndetic. This completes the induction. ∎
The following corollaries of both the Theorem 2.7 and Theorem 2.10 are new.
Corollary 2.11.
Let $n\in\mathbb{N}_{>1},$ $p\in E\left(K(\beta\mathbb{N},\cdot)\right)\cap\overline{\left(K(\beta\mathbb{N},+)\right)}$, and $A\in E_{1}(n,p)$. Then
  1. 1.
    there exists $x,y,\text{ and }z\in\mathbb{N}$ such that
    $\{n^{x},n^{y},n^{z},n^{yz},(n^{x})^{n^{y+z}},(n^{x})^{n^{yz}}\}\subset A,$
    and
  2. 2.
    there exists $a,b,\text{ and }c\in\mathbb{N}$ such that
    $\{n^{a},n^{b},n^{c},n^{b+c},(n^{a})^{n^{b+c}}(n^{a})^{n^{bc}}\}\subset A.$
Proof.
The proof of (1), directly follows from the Corollary 2.8. For the proof of (2), first choose a sequence $\langle x_{t}\rangle_{t=1}^{\infty}$ such that $FS(\langle x_{n}\rangle_{t=1}^{\infty})\subseteq\{m:n^{m}\in A\}\in p,$ and then apply Theorem 2.10. ∎
3   Hindman towers are abundant
As we promised, in this section we improve [22, Theorem 4.1.] in two different directions. As a result, we show that monochromatic Hindman towers are abundant.
Let $(S,\cdot)$ and $(T,\cdot)$ be two discrete semigroups, and $p\in\beta S$ and $q\in\beta T$. Then the tensor product of $p$ and $q$ is defined as
$p\otimes q=\{A\subseteq S\times T:\{x\in S:\{y:(x,y)\in A\}\in q\}\in p\}$
where $x\in S$, $y\in T.$ In their work, Di Nasso and Ragosta [18], considered a groupoid operation $\star$ over $\beta\mathbb{N}$ defined as follows: first, consider the operation $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ defined as $f(n,m)=2^{n}m$. Let $\overline{f}:\beta\left(\mathbb{N}^{2}\right)\rightarrow\beta\mathbb{N}$ be its continuous extension. $\star$ is the restriction of $\overline{f}$ to tensor pairs, namely: for all $p,q\in\beta\mathbb{N}$
$p\star q=\overline{f}(p\otimes q)=2^{p}\odot q.$
The following, which is Theorem 3.1 in [18] is, at the best of our knowledge, the first known result for infinitary exponential configurations in literature, namely Theorem 3.1 in [18].
Theorem 3.1.
[18, Theorem 3.1] (Di Nasso-Ragosta) Let $p\in E\left(K(\beta\mathbb{N},\oplus)\right)$ and $A\in p\star p$. For every $N\in\mathbb{N}$ and for every sequence $\varPhi=\left(f_{n}:n\geq 2\right)$ of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N},$ there exists a sequence $\left(a_{k}\right)_{k\in\mathbb{N}}$ such that
$\mathcal{F}_{k,\varPhi}\left(a_{k}\right)_{k\in\mathbb{N}}=\left\{a_{k}\cdot 2^{\sum_{i=1}^{k-1}\lambda_{i}a_{i}}:k\in\mathbb{N},0\leq\lambda_{1}\leq N,\text{ and }0\leq\lambda_{i}\leq f_{i}\left(a_{i-1}\right)\text{ for }2\leq i\leq k-1\right\}\subset A.$
For each $i\in\mathbb{N}$, choosing $f_{i}$ suitably, we have the exponential Hindman theorem.
Let $N=1$, and choose the function $f_{2}(x)=2^{x}.$ Inductively assume that we have defined the function $f_{n}(x)$ for some $n\in\mathbb{N}_{>1}$. Now define $f_{n+1}(x)=(2^{x})^{f_{n}(x)}$. Now choosing $\varPhi=\left(f_{n}:n\geq 2\right)$, Theorem 3.1 immediately implies the existence of a monochromatic Hindman tower. For details see [18].
For every $p\in\mathbb{P},$ let $deg(p)$ be the degree of the polynomial $p,$ and let $coef(p)=\max\{|c|:c\text{ is a coefficient of }p\}.$ For $n\in\mathbb{N}_{>1},$ define a new operation $\star_{n}$ on $\mathbb{N}$ by $a\star_{n}b=n^{a}b.$ The following theorem is the polynomial extension of Theorem 3.1, established in [22].
Theorem 3.2.
[22, Theorem 4.1.] (Polynomial Di Nasso-Ragosta) Let $p\in E\left(K(\beta\mathbb{N},\oplus)\right)$ and $q\in\beta\mathbb{N}$ be such that for every $N\in\mathbb{N}$, each element of $q$ contains an $IP_{N}$ sets. Let $n\in\mathbb{N}_{>1},$ $A\in p\star_{n}q,$ and $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$. Let $\varPhi=\left(f_{n}:n\geq 2\right)$ be a sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n(>1),x\in\mathbb{N}$, let $F_{n,x}=\{p\in\mathbb{P}:deg(p)\leq f_{n}(x)\text{ and }coef(p)\leq f_{n}(x)\}.$
Under this hypothesis we have a sequence $\left(x_{k}\right)_{k\in\mathbb{N}}$ such that
$\mathcal{PF}^{n}_{k,\varPhi}\left(x_{k}\right)_{k\in\mathbb{N}}=\left\{x_{k}\cdot n^{\sum_{i=1}^{k-1}p_{i}\left(x_{i}\right)}:k\in\mathbb{N},p_{1}\in F_{1},\text{ and }p_{i}\in F_{i,f_{i}\left(x_{i-1}\right)}\text{ for }2\leq i\leq k-1\right\}\subset A.$
Now we find two refinements of Theorem 3.2.
Combined additive and multiplicative patterns in the Hindman towers I
The following theorem is the first improvement of Theorem 3.2. As a corollary we have a monochromatic Hindman tower each of whose terms coming from $MIP_{r}$ sets of a given length.
Theorem 3.3.
Let $n\in\mathbb{N}_{>1},$ $\langle N_{i}\rangle_{i\in\mathbb{N}}$ be a sequence in $\mathbb{N}$, $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, $p\in E(\overline{\left(K(\beta\mathbb{N},+)\right)})$ , $q\in\overline{E((\beta\mathbb{N},+))}$ and $A\in p\star_{n}q$. Let $\varPhi=\left(f_{n}:n\geq 2\right)$ be a sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n\in\mathbb{N}_{>1},G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}=\{p\in\mathbb{P}:deg(p)\leq f_{n}(x)\text{ and }coef(p)\leq f_{n}(x)\text{ for }x=\max G\}.$ Under this hypothesis we have a sequence of finite sets $\langle G_{i}\rangle_{i\in\mathbb{N}}$ such that
  1. (1)
    for each $i\in\mathbb{N},$ $|G_{i}|=N_{i}$, and
  2. (2)
    $\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}=\left\{z_{k}\cdot n^{\sum_{i=1}^{k-1}p_{i}\left(y_{i}\right)}:k\in\mathbb{N},p_{1}\in F_{1},y_{j}\in FS(G_{j})\cup FP(G_{j})\text{ for }1\leq j\leq k-1,\right.$
    $\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,z_{k}\in FS(G_{k})\text{ and }\,p_{i}\in F_{i,f_{i}\left(G_{i-1}\right)}\text{ for }2\leq i\leq k-1\Big{\}}\subset A.$
Proof.
To avoid the complexity in the calculation we prove this theorem up to $k=3.$
The rest of the part can be finished inductively. The technique is verbatim. Let $p\in E\left(K(\beta\mathbb{N},+)\right)$, $q\in E((\beta\mathbb{N},+))$ and $A\in p\star_{n}q=E_{1}(n,p)\cdot q.$ Then $B=\left\{m_{0}:\left\{n_{0}:n^{m_{0}}n_{0}\in A\right\}\in q\right\}\in p.$ Let $m_{1}\in B^{\star}$ and $C_{1}=\left\{n^{\prime}:n^{m_{1}}n^{\prime}\in A\right\}\in q.$ Let $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$. Hence there exists an $IP$ set $FS(\langle n_{1,i}\rangle_{i\in\mathbb{N}})\subseteq C_{1}\in q.$ For each $i\in\mathbb{N},$ let $x_{1,i}=n^{m_{1}}n_{1,i}\in A$. Now $FS(\langle x_{1,i}\rangle_{i\in\mathbb{N}})\subset A.$ By Theorem 2.10, there exists $a_{1}\in B^{\star}$, and $G_{1}\subset FS(\langle x_{1,i}\rangle_{i\in\mathbb{N}})$ such that $|G_{1}|=N_{1},$ and
$\Big{\{}a_{1},a_{1}+p_{1}\left(y_{1}\right):p_{1}\in F_{1},y_{1}\in FS(G_{1})\cup FP(G_{1})\Big{\}}\subset B^{\star}.$
So,
$B_{2}=B^{\star}\cap\left(-a_{1}+B^{\star}\right)\cap\bigcap_{p_{1}\in F_{1}}\bigcap_{y_{1}\in FS(G_{1})\cup FP(G_{1})}-\left(a_{1}+p_{1}\left(y_{1}\right)\right)+B^{\star}\in p.$
Let
$C_{2}=C_{1}\cap\bigcap_{p_{1}\in F_{1}}\bigcap_{y_{1}\in FS(G_{1})\cup FP(G_{1})}\left\{n^{\prime}:n^{a_{1}+p_{1}\left(y_{1}\right)}n^{\prime}\in A\right\}\cap\left\{n^{\prime}:n^{a_{1}}n^{\prime}\in A\right\}\in q.$
Hence there exists an $IP$ set $FS(\langle n_{2,i}\rangle_{i\in\mathbb{N}})\subseteq C_{2}\in q.$ For each $i\in\mathbb{N},$ let $n^{a_{1}}n_{2,i}=x_{2,i}\in A$. Hence $FS(\langle x_{2,i}\rangle_{i\in\mathbb{N}})\subseteq A.$ And for each $y_{1}\in FS(G_{1})\cup FP(G_{1}),$ and $y\in FS(\langle x_{2,i}\rangle_{i\in\mathbb{N}}),$ and $p_{1}\in F_{1}$, we have $yn^{p_{1}\left(y_{1}\right)}\in A.$ As $FS(\langle x_{2,i}\rangle_{i\in\mathbb{N}})\subseteq A$ is an $IP$ set, from Theorem 2.10, we can choose a $G_{2}\subset FS(\langle x_{2,i}\rangle_{i\in\mathbb{N}})$ such that $|G_{2}|=N_{2}$ and
$\left\{a_{2},a_{2}+p_{2}\left(y_{2}\right):p_{2}\in F_{2,f_{2}\left(y_{1}\right)},y_{2}\in FS(G_{2})\cup FP(G_{2})\right\}\subset B_{2}^{\star}.$
Letting $c=a_{1}+a_{2}$, we have
$\left\{c+p_{1}\left(y_{1}\right)+p_{2}\left(y_{2}\right):p_{1}\in F_{1},p_{2}\in F_{2,f_{2}\left(y_{1}\right)},y_{i}\in FS(G_{i})\cup FP(G_{i})\text{ for }1\leq i\leq 2\right\}\subset B^{\star}.$
Let
$C_{3}=C_{1}\cap\bigcap_{p_{1}\in F_{1}}\bigcap_{p_{2}\in F_{2,f_{2}\left(x_{1}\right)}}\bigcap_{1\leq i\leq 2}\bigcap_{y_{i}\in FS(G_{i})\cup FP(G_{i})}\left\{n^{\prime}:n^{c+p_{1}\left(y_{1}\right)+p_{2}\left(y_{2}\right)}n^{\prime}\in A\right\}\in q$
and
$B_{3}^{\star}=B^{\star}\cap\bigcap_{p_{1}\in F_{1}}\bigcap_{p_{2}\in F_{2,f_{2}\left(x_{1}\right)}}\bigcap_{1\leq i\leq 2}\bigcap_{y_{i}\in FS(G_{i})\cup FP(G_{i})}-(c+p_{1}\left(x_{1}\right)+p_{2}\left(x_{2}\right))+B^{\star}\in p.$
Choose an $IP$ set $FS(\langle n_{3,i}\rangle_{i\in\mathbb{N}})\subseteq C_{3}\in q.$ Again for each $i\in\mathbb{N},$ let $n^{c}n_{3,i}=x_{3,i}\in A$. So $FS(\langle x_{3,i}\rangle_{i\in\mathbb{N}})\subseteq A.$ Hence for every $1\leq i\leq 2,$ and $y_{i}\in FS(G_{i})\cup FP(G_{i})$, $z\in FS(\langle x_{3,i}\rangle_{i\in\mathbb{N}})$, $p_{1}\in F_{1},\text{ and }p_{2}\in F_{2,f_{2}\left(x_{1}\right)}.$ we have $zn^{p_{1}\left(y_{1}\right)+p_{2}\left(y_{2}\right)}\in A.$ As $FS(\langle x_{3,i}\rangle_{i\in\mathbb{N}})$ is an $IP$ set, and $B_{3}^{*}$ is piecewise syndetic set, we can use Theorem 2.10 as before to find a set $G_{3}\subset FS(\langle x_{3,i}\rangle_{i\in\mathbb{N}})$ with $|G_{3}|=N_{3}.$
Now iterating this argument we have an infinite sequence of finite subsets $\langle G_{n}\rangle_{n\in\mathbb{N}}$ such that the desired result is true. ∎
Now we have the following corollary which shows that monochromatic Hindman’s towers are abundant.
Corollary 3.4.
Let $\langle N_{i}\rangle_{i}$ be a sequence in $\mathbb{N}.$ Then for every finite coloring of $\mathbb{N},$ there exists a sequence of finite sets $\langle G_{i}\rangle_{i}$ such that $|G_{i}|=N_{i}$ for every $i\in\mathbb{N},$ such that the pattern
$\left\{x_{i_{n}}^{x_{i_{n-1}}^{\cdot^{\cdot^{\cdot^{x_{i_{1}}}}}}}:x_{j}\in FP(G_{j})\,\forall j\in\mathbb{N},\text{ and }1\leq i_{1}<\cdots<i_{n-1}<i_{n},n\in\mathbb{N}\right\}$
is monochromatic.
Proof.
Let $I:\mathbb{N}\rightarrow\mathbb{N}$ be the polynomial defined by for all $x\in\mathbb{N}$, $I(x)=x.$ Let $f_{2}(x)=2^{x},$ and define inductively $f_{n+1}(x)=(2^{x})^{f_{n}(x)}$ for all $n\in\mathbb{N}$. In the above theorem if we choose these sequence of functions $(f_{n})_{n\in\mathbb{N}}$, $F_{n,G}=\{I\}$ for all $G\in\mathcal{P}_{f}(\mathbb{N}),$ then we have a sequence of finite sets $\langle G_{i}\rangle_{i\in\mathbb{N}}$ such that for each $i\in\mathbb{N},$ $G_{i}=\langle n^{x_{i,k}}\rangle_{k=1}^{N_{i}}$ such that the given pattern is monochromatic. ∎
Another application of Theorem 3.3 gives us a new monochromatic pattern. Before that, let us first introduce the notions of $FP_{n}$ and $EFP_{n}$ sets.
Definition 3.5.
For every $n\in\mathbb{N}_{>1},$ and $F\in\mathcal{P}_{f}(\mathbb{N}),$ define
  • $FP_{n}(F)=\{n^{y}:y\in FS(F)\}$,
  • $EFP_{n}(F)=\{n^{y}:y\in FP(F)\}$, and
  • $FEFP_{n}(F)=FP_{n}(F)\cup EFP_{n}(F).$
Clearly $FP_{n}(F)=FP(n^{FS(F)}).$ The following corollary is a straightforward application of Theorem 3.3.
Corollary 3.6.
Let $n,l,r\in\mathbb{N}$, and $\mathbb{N}=\cup_{i=1}^{r}A_{i}$ be a $r$- coloring of $\mathbb{N}$. Then there exist $i\in\{1,2,\ldots,r\}$, and two finite set sets $F,G$ such that
$FP_{n}(F)\cup FP_{n}(G)\cup\{y^{z}:y\in FP_{n}(G),z\in FP_{n}(F)\cup EFP_{n}(F)\}\subset A_{i}.$
Proof.
Choose $G_{1}=F$, and $G_{2}=G$ in Theorem 3.3. ∎
A greedy construction: Let $\mathbb{N}=A\cup A^{c}$ be a two coloring such that $n\in\mathbb{N}_{>1}\cap A$ (resp. in $A^{c}$) implies $n^{n}\in A^{c}$ (resp. in $A$). Then this greedy construction gives us two colors of $\mathbb{N}$ such that none of the colors contain a pattern of the form $\{a,a^{a}\},$ where $a>1.$
But the following corollary of Theorem 3.3 shows that exponential translations of $a^{a}$ by an infinite set lie in the same color with $a.$
Corollary 3.7.
Let $r\in\mathbb{N},$ and $\mathbb{N}=\cup_{i=1}^{r}A_{i}$ be a finite coloring of $\mathbb{N}.$ Then there exist $i,j\in\{1,2,\ldots,r\}$ (possibly $i\neq j$), $a\in\mathbb{N}$, and a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ such that
  1. (1)
    $FS(\langle x_{n}\rangle_{n=1}^{\infty})\subseteq A_{j}$, and
  2. (2)
    $\{a,a^{ya}:y\in FS(\langle x_{n}\rangle_{n=1}^{\infty})\}\subseteq A_{i}$.
Proof.
Let $\mathbb{N}=\cup_{i=1}^{r}B_{i}$ be a $r$ coloring induced by the formulae $m\in B_{i}\iff n^{m}\in A_{i}.$ Let $I:\mathbb{N}\rightarrow\mathbb{N}$ be the polynomial defined by $I(x)=x$ for all $x\in\mathbb{N}.$ Let $q\in E((\beta\mathbb{N},+))$ and $p\in E\left(K(\beta\mathbb{N},+)\right)$ and let $B_{i}\in p*_{n}q$ for some $i\in\{1,2,\ldots,r\}.$ Let $F_{1}=\{I\}.$ Now from the first step of the proof of Theorem 3.3, there exist an element $x$, and a sequence $\langle z_{n}\rangle_{n=1}^{\infty}$ such that $\{z\cdot n^{x}:z\in FS(\langle z_{n}\rangle_{n=1}^{\infty})\}\subseteq B_{i}$. Using pigeonhole principle we have a sub $IP$ set $\langle y_{n}\rangle_{n=1}^{\infty}$ such that $x\cdot FS(\langle y_{n}\rangle_{n=1}^{\infty})\subset FS(\langle z_{n}\rangle_{n=1}^{\infty}).$ Let $n^{x}=a.$ Then $\{a,a^{ya}:y\in FS(\langle y_{n}\rangle_{n=1}^{\infty})\}\subset A_{i}.$
Again from the Hindman theorem [25], $FS(\langle y_{n}\rangle_{n=1}^{\infty})\subseteq\cup_{i=1}^{r}A_{i}$ implies there exists a sub $IP$ set $FS(\langle x_{n}\rangle_{n=1}^{\infty})$ of $FS(\langle y_{n}\rangle_{n=1}^{\infty})$ such that $FS(\langle y_{n}\rangle_{n=1}^{\infty})\subseteq A_{j}.$ Whereas $\{a,a^{ya}:y\in FS(\langle x_{n}\rangle_{n=1}^{\infty})\}\subset A_{i}.$ This completes the proof. ∎
Let $n\in\mathbb{N},$ $\langle N_{i}\rangle_{i\in\mathbb{N}}$ be any sequence in $\mathbb{N}$, $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, and $\varPhi=\left(f_{n}:n\geq 2\right)$ be any sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n(>1)$ and $G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}$ be defined as in the hypothesis of Theorem 3.3. Let $\mathcal{L}_{\mathcal{PF}}$ be the collection of $p\in\beta\mathbb{N}$ such that each $A\in p$ contains an element of the form $\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}$ for some sequence $\langle G_{k}\rangle_{k\in\mathbb{N}}$ of finite sets such that $|G_{k}|=N_{k}$ for each $k\in\mathbb{N}$. Then the following lemma shows that $\mathcal{L}_{\mathcal{PF}}$ is a left ideal of $(\beta\mathbb{N},\cdot).$
Lemma 3.8.
$\mathcal{L}_{\mathcal{PF}}$ is a left ideal of $(\beta\mathbb{N},\cdot).$
Proof.
Let $p\in\mathcal{L}_{\mathcal{PF}}$, and $q\in(\beta\mathbb{N},\cdot).$ Let $A\in q\cdot p.$ Then $B=\{x:x^{-1}A\in p\}\in q.$ Choose $b\in B$, and so $b^{-1}A\in p.$ Let $\langle N_{i}\rangle_{i\in\mathbb{N}}$ be any sequence in $\mathbb{N}$, $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, and $\varPhi=\left(f_{n}:n\geq 2\right)$ be any sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ Then define
  1. 1.
    for each $i\geq 2,$ define $\varphi^{\prime}=(f^{\prime}_{i}:i\geq 2)$ by $f^{\prime}_{i}(x)=f_{i}(bx)$ for all $x\in\mathbb{N}$,
  2. 2.
    for each $i\in\mathbb{N}$, $F^{\prime}_{i}\in\mathcal{P}_{f}(\mathbb{P})$ is defined by $p\in F^{\prime}_{i}$ iff for all $x\in\mathbb{N}$, $p(x)=q(bx)$ for some $q\in F_{i}$ or, $p(x)=r(b^{N_{1}}x)$ for some $r\in F_{i}$.
Now for these new $\varphi^{\prime}$ and $(F^{\prime}_{i}:i\in\mathbb{N})$ there exists a sequence of finite sets $\langle H_{i}\rangle_{i\in\mathbb{N}}$ such that $\mathcal{PF}^{n}_{k,\varPhi}\left(H_{k}\right)_{k\in\mathbb{N}}\subset b^{-1}A.$ Let $\langle G_{i}\rangle_{i\in\mathbb{N}}$ be a sequence of finite sets such that for each $i\in\mathbb{N}$, $G_{i}=\{by:y\in H_{i}\}.$ Then we have $\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset A.$ This completes the proof.
Now from the Ellis’s theorem, we have $\mathcal{L}_{\mathcal{PF}}$ contains a minimal idempotent. The following theorem is an application of Lemma 3.8 which shows that we can mix certain partition regular structures with exponential patterns.
Theorem 3.9.
Let $p\in\mathcal{L}_{\mathcal{PF}}$ be a minimal idempotent and $A\in p$. Then for every linear Rado family $\mathcal{F}$, there exists $F\in\mathcal{F}$ such that for any sequence $\langle N_{i}\rangle_{i\in\mathbb{N}}$ in $\mathbb{N}$, sequence of functions $\varPhi=\left(f_{n}:n\geq 2\right)$ and $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$ we have a sequence of finite sets $(G_{i})_{i\in\mathbb{N}}$ such that
  1. 1.
    for each $i\in\mathbb{N}$, $|G_{i}|=N_{i}$, and
  2. 2.
    $F\cup\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\cup F\cdot\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subseteq A.$
Proof.
As $p\in\mathcal{L}_{\mathcal{PF}}$ is a idempotent, $A\in p$ implies $A^{*}=\{x\in A:x^{-1}A\in p\}$ (from [31, Lemma 4.14]). As $p$ is a minimal ultrafilter, there exists $G\in\mathcal{F}$ such that $G\subset A^{*}.$ Hence $B=A\cap\bigcap_{x\in G}x^{-1}A\in p.$ Again $B\in p$ implies $B^{*}=\{y\in B:y^{-1}B\in p\}\in p.$ Now choose $b\in B^{*}$. Hence the set $C=b^{-1}B\cap B^{*}\in p.$ Let $(G_{k})_{k}$ be a sequence of finite sets such that for each $k\in\mathbb{N}$,
  1. (1)
    $|G_{k}|=N_{k}$, and
  2. (2)
    $\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset C\subseteq A.$
Hence $b\cdot\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset B.$ Denote $F=b\cdot G.$ Then $F\in\mathcal{F}$ and $F\subset A.$ Again we have $b\cdot\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset\bigcap_{x\in G}x^{-1}A$, and this implies $F\cdot\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset A.$
This completes the proof. ∎
The proof of the following corollary is similar to the proof of Corollary 3.4222one has to redefine the sequence of function $\varPhi$ accordingly, so we leave it to the reader.
Corollary 3.10.
[…]
  1. 1.
    for each $i\in\mathbb{N}$, $|G_{i}|=N_{i}$, and
  2. 2.
    $n^{F}\cup\left\{x_{i_{m}}^{x_{i_{m-1}}^{\cdot^{\cdot^{\cdot^{x_{i_{1}}}}}}}:x_{i_{m}}\in FP_{n}(G_{i_{m}}),x_{i_{j}}\in FEFP_{n}(G_{i_{j}})\text{ for }1\leq j\leq m-1\right.$
[…]
    $\left\{x_{i_{m}}^{(a_{k_{m-1}}x_{i_{m-1}})^{\cdot^{\cdot^{\cdot^{(a_{k_{1}}x_{i_{1}})}}}}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!:x_{i_{m}}\in FP_{n}(G_{i_{m}}),x_{i_{j}}\in FEFP_{n}(G_{i_{j}})\text{ for }1\leq j\leq m-1\right.$
[…]
The following example is new and immediately follows from Corollary 3.10. This example shows the existence of monochromatic Hindman tower each of whose terms are coming from a set of arithmetic progressions of a given length.
Example 3.11.
Let $\mathcal{F}$ be the family of finite arithmetic progressions of arbitrary length. Let $p\in\mathcal{L}_{\mathcal{PF}}$ be a minimal idempotent and $A\in E_{1}(n,p)$ where $n\in\mathbb{N}_{>1}$. Then for any $l\in\mathbb{N}$, from Corollary 3.10 we have $F=\{a,a+d,\ldots,a+ld\}\in\mathcal{F}$ such that
$\{n^{a+id}:0\leq i\leq l\}\cup\left\{x_{i_{m}}^{x_{i_{m-1}}^{\cdot^{\cdot^{\cdot^{x_{i_{1}}}}}}}:1\leq i_{1}<\cdots<i_{m-1}<i_{m},m\in\mathbb{N}\right\}\cup$
$\left\{x_{i_{m}}^{((a+j_{m-1}d)x_{i_{m-1}})^{\cdot^{\cdot^{\cdot^{((a+j_{1}d)x_{i_{1}})}}}}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!:1\leq i_{1}<\cdots<i_{m-1}<i_{m},\,and\,j_{1},\ldots,j_{m-1}\in\{0,1,\ldots,l\},m\in\mathbb{N}\right\}\subset A.$
Combined additive and multiplicative patterns in the Hindman tower II
In this section, we find the second improvement of Theorem 3.2. As a corollary, we have monochromatic Hindman towers each of whose terms are coming from a set of arithmetic progressions, geometric progressions.
Theorem 3.12.
Let $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, $n\in\mathbb{N}_{>1}$, and $\langle r_{i}\rangle_{i=1}^{\infty}$, $\langle l_{i}\rangle_{i=1}^{\infty}$ be two sequences in $\mathbb{N}$. Choose $p\in E(\overline{K(\beta\mathbb{N},+)})$, $q\in K(\beta\mathbb{N},\cdot)$, and $A\in p\star_{n}q$. Let $\phi=(f_{n}:n\geq 2)$ be a sequence of functions $f_{n}:\mathbb{N}\longrightarrow\mathbb{N}$. For every $n(>1),G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}=\{p\in\mathbb{P}:deg(p)\leq f_{n}(x)\text{ and }coef(p)\leq f_{n}(x)\text{ for }x=\max G\}.$
Then for each $m\in\mathbb{N}$, there exists elements $x_{m},y_{m},\langle x_{i}^{n}\rangle_{i=1}^{r_{m}}$ in $\mathbb{N}$ such that
  1. 1.
    $R_{m}=FS(\langle x_{i}^{m}\rangle_{i=1}^{r_{m}})\cup\{y_{m},x_{m}+iy_{m}:i=0,\cdots,l_{m}\}\subset A$, and
  2. 2.
    $\mathcal{FG}_{k,\varPhi}^{n}(R_{k})_{k\in\mathbb{N}}=\left\{z_{k}\cdot n^{\sum_{i=1}^{k-1}p_{i}(y_{i})}:k\in\mathbb{N},p_{1}\in F_{1},y_{j}\in R_{j},\text{ for }1\leq j\leq k-1,z_{k}\in R_{k},\right.\\ \left.\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,p_{i}\in F_{i,f_{i}(R_{i-1})}\text{ for }2\leq i\leq k-1\right\}\subset A$.
Proof.
Similar to the proof of Theorem 3.3, we prove this theorem up to the case $k=3.$ The rest of the part can be proved inductively. The technique will be verbatim.
Let $p\in E(\overline{K(\beta\mathbb{N},+)})$, $q\in K(\beta\mathbb{N},\cdot)$ and $A\in p\star_{n}q$. Let $B=\{m_{0}:\{n_{0}:n^{m_{0}}n_{0}\in A\}\in q\}\in p$. Let $m_{1}\in B^{*}$ and $C_{1}=\{n^{\prime}:n^{m_{1}}n^{\prime}\in A\}\in q$. As $C_{1}\in q$, the set $S_{1}=\{x:x^{-1}C_{1}\in q\}$ is syndetic. Given $l_{1},r_{1}\in\mathbb{N}$. Hence there exist $a_{1},d_{1},x_{1}^{1},\cdots,x_{r_{1}}^{1}\in\mathbb{N}$ such that $R_{1}=FS(\langle x_{i}^{1}\rangle_{i=1}^{r_{1}})\cup\{a_{1},d_{1},a_{1}+d_{1},\ldots,a_{1}+l_{1}d_{1}\}\subset S_{1}$. Hence $D_{1}=C_{1}\cap\bigcap_{r\in R_{1}}r^{-1}C_{1}\in q$. For each $p\in F_{1}$ and $r\in R_{1}$, define a new polynomial $p_{r}:\mathbb{Z}\longrightarrow\mathbb{Z}$ by $p_{r}(z)=p(n^{m_{1}}rz)$. Let $G_{1}=\{p_{r}:p\in F_{1},r\in R_{1}\}$ be a new finite set of polynomials from $\mathbb{P}$. As $B^{*}$ is piecewise syndetic, from Theorem 2.3 we have
$E_{1}=\{c:\text{ there exists }b\text{ such that }\{b,b+p(c):p\in G_{1}\}\subset B^{*}\}$
is an $IP_{r_{1}}^{*}$ set for some $r_{1}\in\mathbb{N}$. Hence we have $E_{1}\cap D_{1}\in q$, and let $n_{1}\in E_{1}\cap D_{1}$. As $d_{1}\in R_{1},n_{1}d_{1}\in C_{1}$, and so $n^{m_{1}}n_{1}d_{1}=y_{1}\in A.$ Again,
  1. 1.
    for $i_{1}\in\{0,\ldots,l_{1}\}$, $n^{m_{1}}n_{1}(a_{1}+i_{1}d_{1})=x_{1}+i_{1}y_{1}\in A$ where $x_{1}=n^{m_{1}}n_{1}a_{1}$, and
  2. 2.
    for $i\in\{1,\ldots,r_{1}\}$, $n^{m_{1}}n_{1}FS(\langle x_{i}^{1}\rangle_{i=1}^{r_{1}})=FS(\langle z_{i}^{1}\rangle_{i=1}^{r_{1}})\subset A$ where $z_{i}^{1}=n^{m_{1}}n_{1}x_{i}^{1}$.
Denote $H_{1}=FS(\langle z_{i}^{1}\rangle_{i=1}^{r_{1}})\cup\{y_{1},x_{1}+i_{1}y_{1}:i_{1}=0,\cdots,l_{1}\}$. As $n_{1}\in E_{1},$ choose $b_{1}\in\mathbb{N}$ such that $\{b_{1},b_{1}+p(n_{1}):p\in G_{1}\}\subset B^{*}$. Now from the choice of $G_{1}$, we have $\{b_{1},b_{1}+p(c):p\in F_{1},c\in H_{1}\}\subset B^{*}$. Let
$C_{2}=C_{1}\cap\bigcap_{p\in F_{1},c\in H_{1}}\{n^{\prime}:n^{b_{1}+p(c)}n^{\prime}\in A\}\cap\{n^{\prime}:n^{b_{1}}n^{\prime}\in A\}\in q.$
As $C_{2}$ is piecewise syndetic, the set $S_{2}=\{x:x^{-1}C_{2}\in q\}$ is syndetic. For given $l_{2},r_{2}\in\mathbb{N}$, there exist $a_{2},d_{2},x_{1}^{2},\cdots,x_{r_{2}}^{2}\in\mathbb{N}$ such that $R_{2}=FS(\langle x_{i}^{2}\rangle_{i=1}^{r_{2}})\cup\{a_{2},d_{2},a_{2}+i_{2}d_{2}:i_{2}=0,\cdots,l_{2}\}\subset S_{2}$. Hence $D_{2}=C_{2}\cap\bigcap_{s\in R_{2}}s^{-1}C_{2}\in q$. For each $p\in F_{2,f_{2}(H_{1})}$ and $s\in R_{2}$, define a new polynomial $p_{s}:\mathbb{Z}\longrightarrow\mathbb{Z}$ by $p_{s}(z)=p(n^{b_{1}}sz)$. Let $G_{2}=\{p_{s}:p\in F_{2,f_{2}(H_{1})},s\in R_{2}\}$ be a new finite set of polynomials from $\mathbb{P}$. Let
$B_{2}=B^{*}\cap\bigcap_{p\in F_{1},c\in H_{1}}-(b_{1}+p(c))+B^{*}\cap(-b_{1}+B^{*})\in p.$
Now
$E_{2}=\{d:\text{there exists }a\text{ such that }\{a,a+p(d):p\in G_{2}\}\subset B_{2}^{*}\}$
is an $IP_{t}^{*}$ set for some $t\in\mathbb{N}$. Choose $n_{2}\in E_{2}\cap D_{2}$. For $i\in\{1,\cdots,r_{2}\},$ let $n^{b_{1}}n_{2}x_{i}^{2}=z_{i}^{2}$, $n^{b_{1}}n_{2}a_{2}=x_{2}$, and $n^{b_{1}}n_{2}d_{2}=y_{2}$. Clearly $H_{2}=FS(\langle z_{i}^{2}\rangle_{i=1}^{r_{2}})\cup\{y_{2},x_{2}+i_{2}y_{2}:0\leq i_{2}\leq l_{2}\}\subset A$, and $\{h_{2}n^{p(h_{1})}:h_{2}\in H_{2},h_{1}\in H_{1},p\in F_{1}\}\subset A$.
Again choose $b_{2}\in\mathbb{N}$ such that $\{b_{2},b_{2}+p(n_{2}):p\in G_{2}\}\subset B_{2}^{*}$. Hence, for all $p_{1}\in F_{1},p_{2}\in F_{2,f_{2}(H_{1})},h_{1}\in H_{1},$ and $h_{2}\in H_{2}$, we have $b_{1}+b_{2}+p_{1}(h_{1})+p_{2}(h_{2})\in B^{*}$. Now proceeding iteratively, we have our desired result. ∎
Before we apply Theorem 3.12, let us introduce the the following notion of large set.
Definition 3.13.
For given $m,n\in\mathbb{N}$, a $FGP_{m,n}$ set is a set of the form $FP(\langle x_{i}\rangle_{i=1}^{m})\cup\{d\}\cup\{ad^{i}:i\in\{0,1,\ldots,n\}\}$ for some $a,d,\langle x_{i}\rangle_{i=1}^{m}$ in $\mathbb{N}$. That is these sets consist of an multiplicative $IP_{m}$ set and a set of geometric progressions of length $n.$
Let $n\in\mathbb{N}_{>1},$ $\langle r_{i}\rangle_{i=1}^{\infty}$, $\langle l_{i}\rangle_{i=1}^{\infty}$ be two sequences in $\mathbb{N}$. Let $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, and $\varPhi=\left(f_{n}:n\geq 2\right)$ be any sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n\in\mathbb{N}_{>1},$ and $G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}$ be defined as in the hypothesis of Theorem 3.12. Let $\mathcal{R}_{\mathcal{F}\mathcal{G}}$ be the collection of $p\in\beta\mathbb{N}$ such that each $A\in p$ contains a set of the form $\mathcal{FG}_{k,\varPhi}^{n}(R_{k})_{k\in\mathbb{N}}$ for some sequence of finite sets $(R_{k})_{k\in\mathbb{N}}$, where each $R_{k}$ is a $FGP_{r_{k},l_{k}}$ set. Then the following lemma shows that $\mathcal{R}_{\mathcal{FG}}$ is a left ideal of $(\beta\mathbb{N},\cdot).$
Lemma 3.14.
$\mathcal{R}_{\mathcal{FG}}$ is a left ideal of $(\beta\mathbb{N},\cdot).$
Proof.
The proof is similar to the proof of Lemma 3.8, and so we leave it to the reader. ∎
The following theorem is an analog of Theorem 3.9, and shows that we can mix exponential patterns with members from a given linear Rado family.
Theorem 3.15.
Let $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$, $p\in\mathcal{R}_{\mathcal{FG}}$ be a minimal idempotent and $A\in p$. Then for every linear Rado family $\mathcal{F}$, there exists $F\in\mathcal{F}$ such that for any two sequences $\langle r_{i}\rangle_{i=1}^{\infty}$, $\langle l_{i}\rangle_{i=1}^{\infty}$ in $\mathbb{N}$, and sequence of functions $\varPhi=\left(f_{n}:n\geq 2\right)$ we have a sequence of finite sets $(R_{k})_{k\in N}$ such that
  1. 1.
    each $R_{k}$ is a $FGP_{r_{k},l_{k}}$ set, and
  2. 2.
    $F\cup\mathcal{FG}_{k,\varPhi}^{n}(R_{k})_{k\in\mathbb{N}}\cup F\cdot\mathcal{FG}_{k,\varPhi}^{n}(R_{k})_{k\in\mathbb{N}}\subset A.$
Proof.
The proof is similar to the proof of Theorem 3.9. ∎
The following Corollary of the above theorem immediately extends the Example 3.11, which shows that monochromatic Hindman towers exist each of whose members coming from a set of arithmetic progressions, geo-metric progressions of a given length.
Corollary 3.16.
Let $l\in\mathbb{N}$, $\langle r_{i}\rangle_{i=1}^{\infty}$ and $\langle l_{i}\rangle_{i=1}^{\infty}$ be two sequences in $\mathbb{N}$. Then for any finite partition of $\mathbb{N}$, there exists a sequence of finite sets $\langle R_{i}\rangle_{i=1}^{\infty}$ such that each $R_{i}$ is a $FGP_{r_{i},l_{i}}$ set, and a set $F=\{a,a+d,\ldots,a+ld\}$ such that
$\{n^{a+id}:0\leq i\leq l\}\cup\left\{x_{i_{n}}^{x_{i_{n-1}}^{\cdot^{\cdot^{\cdot^{x_{i_{1}}}}}}}:x_{j}\in R_{j}\,\forall j\in\mathbb{N},\text{ and }1<i_{1}<\cdots<i_{n-1}<i_{n},n\in\mathbb{N}\right\}\cup$
$\left\{x_{i_{m}}^{((a+j_{m-1}d)x_{i_{m-1}})^{\cdot^{\cdot^{\cdot^{((a+j_{1}d)x_{i_{1}})}}}}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!:x_{j}\in R_{j}\,\forall j\in\mathbb{N},1<i_{1}<\cdots<i_{m},\,and\,j_{1},\ldots,j_{m-1}\in\{0,1,\ldots,l\},m\in\mathbb{N}\right\}$
is monochromatic.
Proof.
The proof is similar to the proof of Corollary 3.4. ∎
The following lemma is a strengthening of [22, Lemma 5.22.], plays a crucial role to show that multiplicative piecewise syndetic sets are exponentially rich.
Lemma 3.17.
Let $n\in\mathbb{N}_{>1}$, and $L^{a}$ and $L^{m}$ be both left ideals of $(\beta\mathbb{N},+)$ and $(\beta\mathbb{N},\cdot).$ Then there exist $p\in L^{a}$ and $q\in L^{m}$ such that $p\star_{n}q=q.$
Infact we can choose $p$ to be an idempotent ultrafilter.
Proof.
In the proof of [22, Lemma 5.22.], we can choose an idempotent $p\in L^{a}$ using [32, Lemma 2.1]. The rest of the proof is the same. ∎
Now we use the above lemma to prove the following corollary of Theorem 3.12 to show that every multiplicative piecewise syndetic set witnesses the conclusion of Theorem 3.12.
Corollary 3.18.
Every multiplicative piecewise syndetic set witness the conclusion of Theorem 3.12. In other words
$\overline{K\left(\beta\mathbb{N},\cdot\right)}\subseteq\mathcal{R}_{\mathcal{FG}}.$
Proof.
Let $L$ be any minimal left ideal of $(\beta\mathbb{N},\cdot)$. Then from Lemma 3.17, there exists $q\in L$, and $p\in E(\overline{K(\beta\mathbb{N},+)})$ such that $p\star_{n}q=q.$ Then from Theorem 3.12, $q\in\mathcal{R}_{\mathcal{FG}}$. Now from Lemma 3.14, $L=\beta\mathbb{N}\cdot q\subset\mathcal{R}_{\mathcal{FG}}$. As $L$ is arbitrary, we have $K\left(\beta\mathbb{N},\cdot\right)=\bigcup\{L:L\text{ is a minimal left idealof }(\beta\mathbb{N},\cdot)\}\subseteq\mathcal{R}_{\mathcal{FG}}.$ Now choose any $p\in\overline{K\left(\beta\mathbb{N},\cdot\right)},$ and $A\in p.$ Now choose $q\in K\left(\beta\mathbb{N},\cdot\right)$ be such that $A\in q.$ But $q\in\mathcal{R}_{\mathcal{FG}}$ and so the set $A$ satisfies the conclusion of the Theorem 3.12. As $A\in p$ is arbitrarily chosen, we have $p\in\mathcal{R}_{\mathcal{FG}}$ This completes the proof. ∎
As a consequence we have for every $n(>1)\in\mathbb{N}$, and $p\in\overline{K\left(\beta\mathbb{N},\cdot\right)}$, $E_{1}(n,p)$ witness the pattern of Corollary 3.16.
Combined additive multiplicative and exponential Hindman theorem
In [8, Theorem 2.4.], V. Bergelson and N. Hindman proved the following combined extension of the additive and multiplicative Hindman theorem. The proof uses those ultrafilters each of whose members are additive and multiplicative central sets.
Theorem 3.19.
[8, Theorem 2.4.] If $\mathbb{N}$ is finitely colored, then there exist two sequences $\langle x_{n}\rangle_{n}$ and $\langle y_{n}\rangle_{n}$ such that $FS(\langle x_{n}\rangle_{n})\cup FP(\langle y_{n}\rangle_{n})$ is in the same color.
In [22, Theorem 4.5.], the authors extended the above theorem for exponential patterns. In the following theorem we show that for any finite coloring of $\mathbb{N}$, there exists a color witnessing the conclusion of Theorem 3.3, 3.12, and 3.19. This simultaneously extends Theorem 3.3, 3.12 and [8, Theorem 2.4.].
Theorem 3.20.
Let $r,n(>1)\in\mathbb{N}$, $\langle N_{i}\rangle_{i\in\mathbb{N}}$, $\langle r_{i}\rangle_{i=1}^{\infty}$ and $\langle l_{i}\rangle_{i=1}^{\infty}$ be three sequences in $\mathbb{N}$, and $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$. Let $\varPhi=\left(f_{n}:n\geq 2\right)$ be a sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n\in\mathbb{N}_{>1},G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}=\{p\in\mathbb{P}:deg(p)\leq f_{n}(x)\text{ and }coef(p)\leq f_{n}(x)\text{ for }x=\max G\}.$ Now for every finite coloring $\mathbb{N}=\cup_{i=1}^{r}C_{i}$, there exists $j\in\{1,2,\ldots,r\},$ two sequences of finite sets $\langle G_{i}\rangle_{i\in\mathbb{N}}$, and $\langle R_{i}\rangle_{i\in\mathbb{N}}$, and two sequences $\langle x_{n}\rangle_{n\in\mathbb{N}}$, and $\langle y_{n}\rangle_{n\in\mathbb{N}}$ such that
  1. 1.
    for each $i\in\mathbb{N},$ $|G_{i}|=N_{i}$,
  2. 2.
    for each $i\in\mathbb{N},$ $R_{i}$ is a $FGP_{r_{i},l_{i}}$ set, and
  3. 3.
    $FS(\langle x_{n}\rangle_{n\in\mathbb{N}})\cup FP(\langle y_{n}\rangle_{n\in\mathbb{N}})\cup\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\cup\mathcal{FG}^{n}_{k,\varPhi}\left(R_{k}\right)_{k\in\mathbb{N}}\subseteq C_{j}.$
Proof.
Let $p\in E(\beta\mathbb{N},+)$ and $q\in\overline{E(\beta\mathbb{N},+)}$. Let $r=p\star_{n}q$ and $L_{r}=\beta\mathbb{N}\cdot r$. Then by Lemma 3.8, $L_{r}$ is the left ideal of $(\beta\mathbb{N},\cdot)$, and so contains a minimal left ideal say $L.$ From Lemma 3.8, each member of $L$ witness the conclusion of Theorem 3.3. Let $t\in L$ be a minimal idempotent and let $t=s\cdot r=(s\cdot E_{1}(n,p))\cdot q.$
Assume that $C_{j}\in t.$ Now from Lemma 3.8, there exists a sequence of finite sets $\left(G_{k}\right)_{k\in\mathbb{N}}$ satisfying for each $i\in\mathbb{N},$ $|G_{i}|=N_{i}$, and $\mathcal{PF}^{n}_{k,\varPhi}\left(G_{k}\right)_{k\in\mathbb{N}}\subset C_{j}.$ Again $t\in K(\beta\mathbb{N},\cdot),$ hence from Lemma 3.18, there exists a sequence of finite sets $\left(R_{k}\right)_{k\in\mathbb{N}}$ such that for each $i\in\mathbb{N},$ $R_{i}$ is a $FGP_{r_{i},l_{i}}$ set, and $\mathcal{FG}^{n}_{k,\varPhi}\left(R_{k}\right)_{k\in\mathbb{N}}\subseteq C_{j}.$ As $t$ is an idempotent of $(\beta\mathbb{N},\cdot)$, there exists a sequence $\langle y_{n}\rangle_{n\in\mathbb{N}}$ such that $FP(\langle y_{n}\rangle_{n\in\mathbb{N}})\subseteq C_{j}.$ As $q\in\overline{E(\beta\mathbb{N},+)}$, and $\overline{E(\beta\mathbb{N},+)}$ is a left ideal of $(\beta\mathbb{N},\cdot)$, we have $t\in\overline{E(\beta\mathbb{N},+)}.$ Hence there exists a sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ such that $FS(\langle x_{n}\rangle_{n\in\mathbb{N}})\subseteq C_{j}.$
This completes the proof. ∎
4   Open questions
In this section, we summarize some of the questions that we could not solve, as well as some possible directions we believe could be interesting to pursue in the future. The first question is regarding the refinement of Corollary 2.11. In that corollary we are unable to insert both additive and multiplicative structure in the exponent of $n$. So one may expect the following joint extension of (1), and (2) of Corollary 2.11 is possible.
Question 4.1.
Let $n\in\mathbb{N}_{>1}$, and $\mathbb{N}$ be finitely colored. Does there exists $x,y,\text{ and }z\in\mathbb{N}$ such that
$\{n^{x},n^{y},n^{z},n^{yz},n^{y+z},(n^{x})^{n^{y+z}},(n^{x})^{n^{yz}}\}$
is monochromatic?
But if the above question has an affirmative answer, then this would imply that the pattern $\{x,y,x+y,xy\}$ is partition regular. But this is a long standing open problem. For recent development in this area, one can look into the articles [1, 2, 15, 16, 33, 34]. But we expect a relatively weak version of the Question 4.2 is true. So we make the following conjecture.
Conjecture 4.2.
Let $n\in\mathbb{N}_{>1}$, and $\mathbb{N}$ be finitely colored. Then there exists $x,y,\text{ and }z\in\mathbb{N}$ such that
$\{n^{x},n^{yz},n^{y+z},(n^{x})^{n^{y+z}},(n^{x})^{n^{yz}}\}$
is monochromatic. In other words the pattern $\{x,y+z,yz,xn^{y+z},xn^{yz}\}$ is partition regular.
The second question is regarding Theorem 3.3. In Theorem 3.3, we proved that for each $k\in\mathbb{N}$, $z_{k}\in FS(G_{k})$. But there is a drawback when we compare it with Theorem 3.12. Because in the later case, we were able to show $z_{k}\in R_{k}$. This immediately requires a refinement of Theorem 3.3, asking whether $z_{k}\in FS(G_{k})\cup FP(G_{k})?$ So we leave the following question open.
Question 4.3.
Let $n\in\mathbb{N}_{>1},$ $\langle N_{i}\rangle_{i\in\mathbb{N}}$ be a sequence in $\mathbb{N}$, $F_{1}\in\mathcal{P}_{f}(\mathbb{P})$. Let $\varPhi=\left(f_{n}:n\geq 2\right)$ be a sequence of functions $f_{n}:\mathbb{N}\rightarrow\mathbb{N}.$ For every $n(>1),G\in\mathcal{P}_{f}(\mathbb{N})$, let $F_{n,G}=\{p\in\mathbb{P}:deg(p)\leq f_{n}(x)\text{ and }coef(p)\leq f_{n}(x)\text{ for }x=\max G\}.$ Now for every finite coloring of $\mathbb{N}$, does there exists a sequence of finite sets $\langle G_{i}\rangle_{i\in\mathbb{N}}$ such that
  1. 1.
    for each $i\in\mathbb{N},$ $|G_{i}|=N_{i}$, and
  2. 2.
    $\left\{y_{k}\cdot n^{\sum_{i=1}^{k-1}p_{i}\left(y_{i}\right)}:k\in\mathbb{N},p_{1}\in F_{1},y_{j}\in FS(G_{j})\cup FP(G_{j})\text{ for }1\leq j\leq k,\right.$
[…]
is monochromatic?
While proving Corollary 3.18, we used the property: for any $q\in K(\beta\mathbb{N},\cdot)$, and some suitably chosen $p\in\beta\mathbb{N}$, $p\star_{n}q$ witness the conclusion of Theorem 3.12. But we could not prove Theorem 3.3 for any $q\in K(\beta\mathbb{N},\cdot)$. That’s why the technique of the proof of Corollary 3.18 can’t be adapted for the case of $\mathcal{L}_{\mathcal{PF}}$. So we leave the following question open.
Question 4.4.
Does every multiplicative piecewise syndetic set satisfies the conclusion of 3.3? In other words does $\overline{K\left(\beta\mathbb{N},\cdot\right)}\subseteq\mathcal{L}_{\mathcal{PF}}?$
Acknowledgement:
The author of this paper is supported by NBHM postdoctoral fellowship with reference no:
0204/27/(27)/2023/R & D-II/11927.